Two years ago Mark was three times as old as Mia. This year he is twice as old. Is that possible? If so, how old is MArk? How old is Mia?
Found 2 solutions by Alan3354, lwsshak3:
Answer by Alan3354(69443)

Two years ago Mark was three times as old as Mia. This year he is twice as old. Is that possible? If so, how old is MArk? How old is Mia?
-------------------
K = Mark's age now
M = Mia's age now
---
K-2 = 3(M-2)
K = 2M
---
Sub for K
2M-2 = 3(M-2) = 3M-6
Add 6
2M+4 = 3M
Subtract 2M
4 = M
M = 4

Answer by lwsshak3(11628)

Two years ago Mark was three times as old as Mia. This year he is twice as old. Is that possible? If so, how old is MArk? How old is Mia?
let x=Mia's age this year
2x==Mark's age this year
2 yrs ago:
x-2=mia's age
2x-2=Mark's age
2x-2=3(x-2)
2x-2=3x-6
x=4
Mia's age=4
Mark's age=8